Ticket Number: HEA-NOC/20080107-1 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080107 10:24 UTC Problem Start: 20080109 07:00 UTC
Ticket Update: 20080109 20:56 UTC Problem End: 20080109 09:00 UTC
Site/line: ABILENE, An Cheim, BCFE, Blacknight, Boston College, CAO,
CHIU, CIT, CMOD, CSM, Canarie, Carriganore, Cluain Mhuire,
College St., Crumlin VEC, DCU, DIAS, DIT, DKIT, Digital
Depot, EPA, Fulbright, GDA, GEANT, GMIT, GRID-Ireland, GSI
Colo, GTI, HEA, HEA-NOC, HETAC, HMRC, HRB, HSE Colo,
Hermes, IADT, ICHEC, IEDR, ILUG, INEX, IRCSET, IT Athlone,
IT Blanchardstown, IT Carlow, IT Cork, IT Dundalk, IT
Sligo, IT Tralee, ITTallaght, ITnet, IUA, JANET, Kilester
VEC, LIT, LYIT, Liberties VEC, MLE, Mater Dei, NCAD, NCIRL,
NCRI, NIRAN, NLI, NQAI, NUI, NUIG, NUIG Clinical Sciences,
NUIM, NYSERNET, Oireachtas, Qualifax, RCSI, RIA, RIFTFS,
RTE, SAD, SINET, STARTAP, SVUH, Schools Network, Sonas
Innovation, St. Angela's, TCD, TCD Tallaght Hospital,
TWAREN, Teagasc, Tipp Inst, Tony Ryan Academy, UCC, UCD,
UL, UL Mary Immaculate, UL Outreach Centre, WIT, WebSat,
Bluenet 7600s require that the interfaces connecting the head-end switches
be changed from routing mode to switched mode.
This is scheduled to take place between 7:00am and 9:00am, Wednesday 9th
The remaining 7600 PE devices will have their access ports changed from
routed to switched between 7:00am and 9:00am, Wednesday 16th Jan 2008.
Each site will experience a momentary outage for all connected sites as the
change is made.
20080107 10:18 GAR Created ticket to track work.
20080109 07:18 GAR No alarms or network issues to preclude change going
ahead. Snapshotting current configuration on PE nodes and starting change.
Initial nodes to be updated shall be DCU and WIT PE nodes.
20080109 08:55 GAR Problem found with dot1Q tagging when change made to
WIT-PE node. Carlow IT Commodity IP service unavailable due to Layer 2
tunnel being unavailable. Reverting change.
20080109 09:20 GAR Fault with ITC Commodity IP traffic link isolated and
20080109 20:52 GAR Modification of remaining 7600 PEs is scheduled to take
place between 7:00am and 9:00am, Wednesday 16th Jan 2008.
This ticket can be monitored at http://www.hea.net/tickets/20080107-1
HEAnet Limited. Registered in Ireland, No. 275301.